OpenStudy (anonymous):

Luis knows that his pool holds 3,050 gallons and that his water pump will pump the water out at a rate of 9 gallons per minute. How long will it take Luis to have an empty pool? y = -9x - 3050 9x – y = -3050 y = 9x + 3050 9x + y = 3050

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

i think it might be A ?

OpenStudy (anonymous):

@Springtrap550

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

or d ?

jimthompson5910 (jim_thompson5910):

`his water pump will pump the water out at a rate of 9 gallons per minute` ---> the slope is -9 `his pool holds 3,050 gallons` this is the starting amount so the y intercept is b = 3050

rebeccaxhawaii (rebeccaxhawaii):

its d

jimthompson5910 (jim_thompson5910):

m = -9 b = 3050 y = mx+b turns into y = -9x + 3050

jimthompson5910 (jim_thompson5910):

if you were to rearrange y = -9x + 3050 into the form Ax + By = C, you'll get 9x + y = 3050

OpenStudy (anonymous):

thank you